An aircraft can fly at a velocity of 280 km/h with no wind interference. It must fly to an airport 275 km/h of its current position. A north-westerly wind with a velocity of 72 km /h is blowing. 
Calculate the direction ij which the aircraft must fly to get to the airport.
1
Expert's answer
2021-05-16T18:00:20-0400
"V'^2=V^2+u^2-2Vu\cos{\theta}\\
275^2=280^2+72^2+2(280)(72)\cos{\theta}\\\theta=101\degree"